Write an equivalent expression using negative exponents 6 over y to the third power
Answer by Theo(3051)

6/y^3 is the same as 6*y^(-3)
in general:
x^(-a) = 1 / x^a
1 / x^(-1) = x^a
the variable becomes it's reciprocal and the exponent changes it's sign.
some examples:
5^2 = 25
5^(-2) = 1/25
1/5^2 = 1/25
1/5^(-2) = 25
you can use your calculator to confirm.